OpenStudy (anonymous):

Solve n^2 - 3n = 0 by factoring. Select all possible answers. Infinite solutions No solution Others that aren't listed here. 3 0 I got Infinite solutions and Others that aren't listed. Am I right?

OpenStudy (mathstudent55):

Didn't I just respond to this question?

OpenStudy (mathstudent55):

Factor n out of the left side.

OpenStudy (haseeb96):

the roots are 0,3 but the correct answer is others that are not listed is the

OpenStudy (anonymous):

Yes but I was given the wrong answe.r..

OpenStudy (mathstudent55):

@Haseeb96 You need to read the problem more carefully.

OpenStudy (mathstudent55):

Let's do together what I wrote above. First, we factor out n. n^2 - 3n = 0 n(n - 3) = 0 @Keridavis ok so far?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

Now we set each factor equal to zero. n = 0 or n - 3 = 0 OK?

OpenStudy (haseeb96):

@Keridavis u have asked a lot of questions of factorization but u dont know how to solve it why?

OpenStudy (mathstudent55):

Now we solve the right equation, and we get n = 0 or n = 3 The answer is choices D and E.